OpenStudy (anonymous):

Someone help please! solve by graphing -5x^2-20x-15=0

OpenStudy (anonymous):

it's a parabola that opens down, the solutions are x = -3 and x = - 1 (these are the x intercepts). To find the vertex we use the vertex formula: h=-b/2a and k=f(h); the vertex is (h,k); so in this case it's the point (-2, 5). With these 3 points you could graph it

OpenStudy (anonymous):

Oh ok so its like this?

OpenStudy (anonymous):

It wont let me draw... But its like a U shape but upside down?

OpenStudy (anonymous):

yes an U shape upside down that starts at (-2,5) and intercepts x at -3 and -1.

OpenStudy (anonymous):

thank you